The perimeter of a rectangle is 34 feet and its area is 60 square feet. Find the length and width of the rectangle.
Answer by checkley75(3666)

xy=60
x=60/y
2x+2y=34
2(60/y)+2y=34
120/y+2y=34
(120+2y^2)/y=34
120+2y^2=34y
2y^2-34y+120=0
2(y^2-17y+60)=0
2(y-12)(y-5)=0
y-12=0
y=12 answer. the x=60/12=5
y-5=0
y=5 answer. the y=60/5=12
proof
5*12=60
60=60
2*5+2*12=34
10+24=34
34=34